### v1

@@ Abstract
The accurate prediction of phase diagrams is of central importance for both the fundamental understanding of materials as well as for technological applications in material sciences. However, the computational prediction of the relative stability between phases based on their free energy is a daunting task, as traditional free energy estimators require a large amount of simulation data to obtain uncorrelated equilibrium samples over a grid of thermodynamic states. In this work, we develop deep generative machine learning models for entire phase diagrams, employing normalizing flows conditioned on the thermodynamic states, e.g., temperature and pressure, that they map to. By training a single normalizing flow to transform the equilibrium distribution sampled at only one reference thermodynamic state to a wide range of target temperatures and pressures, we can efficiently generate equilibrium samples across the entire phase diagram. Using a permutation-equivariant architecture allows us, thereby, to treat solid and liquid phases on the same footing. We demonstrate our approach by predicting the solid-liquid coexistence line for a Lennard-Jones system in excellent agreement with state-of-the-art free energy methods while significantly reducing the number of energy evaluations needed.
@@ 1 Introduction
In recent years, machine learning (ML) models have found broad use in molecular simulations. ML interatomic potentials have been shown to strongly reduce the cost of ab initio MD simulations while retaining an accuracy comparable to density functional theory or other quantum chemistry methods [8, 9, 10]. Lately, these potentials have also been applied in the calculation of phase diagrams [11, 12] providing faster energy evaluations while the sampling challenge remains. Complementary, deep generative models are being developed to generate uncorrelated samples of molecular structures in one go, with the aim of overcoming the sampling problem typical of MD and MC. In particular, the Boltzmann Generator approach [13] leverages a combination of highly expressive invertible coordinate transformations, such as normalizing flows [14, 15], and statistical mechanics methods such as importance sampling or Metropolis-Hastings Monte Carlo, to generate independent identically distributed samples of the targeted equilibrium distribution. This approach has emerged both in the molecular sciences [13] as well as in lattice field theory [16, 17]. Within the TFEP-based learned free energy perturbation (LFEP), using flows enables to learn flexible mappings for systems where physical intuition is insufficient [18]. This approach has successfully been applied to sample small solvation systems [19], monoatomic solids [20, 21, 22], and small molecules with complex quantum-mechanical potentials based on cheaper potentials [23] or at different temperatures to overcome slow modes [24]. However, while LFEP offers an elegant way to learn a mapping between two specific thermodynamic states, it cannot be used efficiently for the calculation of phase diagrams since a new flow model needs to be trained for each thermodynamic state.
@@
In this work, we generalize the LFEP and Boltzmann Generator approaches by employing normalizing flows conditioned on the target thermodynamic states. This conditional flow is capable of mapping MD samples from only one reference thermodynamic state across the entire $(T,P)$-range of the phase diagram with a single trained model. By training to match the equilibrium distributions at all thermodynamic states and appropriately reweighting the generated samples, free energy differences can be very efficiently computed for any temperature and pressure. In this way, the coexistence line between phases can be mapped out with only minimal prior knowledge of its location, while avoiding the need to conduct numerous MD simulations across a grid of thermodynamic states. Using a permutation-equivariant architecture [18], our approach further treats ordered and disordered phases on the same footing. We demonstrate our developed framework by determining the coexistence line between solid and liquid phases of a Lennard-Jones (LJ) system consisting of 180 interacting particles, achieving highly accurate free energy estimates for both phases over the entire range of temperatures and pressures.
Figure 1: Workflow conditional flow.a, The prior equilibrium distribution $q_{{\bf c}_{0}}$ at a reference thermodynamic state ${\bf c}_{0}=(T_{0},P_{0})$ is sampled with MD. A flow conditioned on the thermodynamic variables transforms this distribution into $q^{\prime}_{0i}$ which approximate the target equilibrium distributions $q_{{\bf c}_{i}}$ at different thermodynamic states ${\bf c}_{i}$. The flow is trained by minimizing the conditional loss function $\mathcal{L(\theta)}$, given as expectation value over the conditional value ${\bf c}\sim p_{\bf c}$ of the single-point losses. b, Schematic of the conditional transformation of one sample $f({\bf x},V)=({\bf x}^{\prime},V^{\prime})$ consisting of an atomic configuration ${\bf x}$ and a volume parameter $V$. The transformation of the box parameter is a simple isotropic scaling operation ($L^{\prime}_{i}/L^{\prime}_{j}\!=\!L_{i}/L_{j}$) learned by an MLP conditioned on $(T,P)$. The configuration is scaled to fractional coordinates ${\bf s}$ and then transformed by coupling layers conditioned on $T$, $P$, and $V^{\prime}$, which transform a subset of cartesian coordinates of all particles in each coupling layer (see Flow architecture for details), before it is scaled back to real coordinates.
2 Results
2.1 Conditioning normalizing flows on temperature and pressure
Normalizing flows [14, 15] are a class of invertible deep generative models that, given samples from a prior distribution, generate samples from a new distribution. In contrast to other generative models, the use of invertible transformations allows to evaluate the generated distribution using the change-of-variables technique. By minimizing the Kullback-Leibler (KL) divergence between the generated and target distribution, flows can be trained to produce samples from this target distribution. Flows can further be trained in a conditional setting which allows the modeling of conditional target distributions. While originally developed in the context of image generation tasks [25, 26], the idea of conditioning flows has recently been applied to the sampling of rare events and to lattice field theories [27, 28]. Within LFEP, parametrizing the configurational mapping as a normalizing flow allows to sample the target state given samples from the prior state together with the subsequent evaluation of the free energy difference between the two states.
The key idea behind our approach is to compute free energy differences over a wide range of thermodynamic states without the need to run MD simulations at each state point. This is achieved by training a normalizing flow conditioned on temperature and pressure, which learns the mapping from a prior distribution at a reference thermodynamic state to a whole family of target distributions across thermodynamic space. Figure 1 a) illustrates the workflow of our approach. We first sample the prior distribution at the reference state in the $NPT$ ensemble with MD and save samples consisting of the atomic configuration ${\bf x}$ and the box volume $V$. In this work, we consider rectangular boxes and only allow for isotropic scaling, but the proposed approach is completely general and can easily be extended to fully triclinic simulation boxes. Samples $({\bf x},V)$ from the prior distribution $q_{{\bf c}_{0}}$ of the reference state ${\bf c}_{0}=(T_{0},P_{0})$ with temperature $T_{0}$ and pressure $P_{0}$ are transformed by the flow to a generated configuration and volume in order to approximate equilibrium distributions at multiple different thermodynamic conditions. While the volume is updated using an affine transformation conditioned on the thermodynamic state, the configuration is first scaled to fractional coordinates and then transformed by a flow conditioned on the updated volume as well as on the thermodynamic state. For both transformations, the conditioning is achieved by feeding the thermodynamic state as an additional input to the model during training (Fig. 1 b[25].
To optimize the flow parameters, we employ conditional energy-based training, avoiding the need to sample the target distributions. In particular, the average of the KL divergences between the generated and the target distributions conditioned on each target state over the entire range of thermodynamic states is evaluated, yielding the following loss function (see 4.2 for a derivation):
$\mathcal{L}(\theta)=-\mathbb{E}_{(T,P)\sim p_{\bf c}}\mathbb{E}_{({\bf x},V)\sim q_{{\bf c}_{0}}}\bigl{[}-\beta_{0}(U({\bf x})$ $+P_{0}V)$ (1)
$+\beta(U({\bf x}^{\prime})$ $+PV^{\prime})$
[…]
Here, $U({\bf x})$ and $\beta=1/k_{B}T$ denote the potential energy and inverse temperature, respectively. The conditioning states ${\bf c}=(T,P)$ are drawn from the distribution $p_{\bf c}$, and $({\bf x}^{\prime},V^{\prime})$ is the sample generated by the conditional flow $f({\bf x},V|T,P)$ with parameters $\theta$ and Jacobian $J_{f}$. After training, equilibrium expectation values of an observable $O({\bf x},V)$ can be computed as
$\mathbb{E}_{({\bf x},V)\sim q_{\bf c}}\,\left[O({\bf x},V)\right]\approx\frac{\mathbb{E}_{({\bf x},V)\sim q_{{\bf c}_{0}}}\left[w({\bf x},V|{\bf c})O({\bf x}^{\prime},V^{\prime})\right]}{\mathbb{E}_{({\bf x},V)\sim q_{{\bf c}_{0}}}\left[w({\bf x},V|{\bf c})\right]}$ (2)
using the conditional importance weights $w({\bf x},V|{\bf c})$ (Eq. (10)). Compared to temperature steerable flows [29], the proposed workflow is more general and allows for more flexible models, as no requirements need to be placed on the architecture or the prior.
As condensed phase systems can exist in the liquid as well as in the solid phase, it is crucial to treat both phases on the same footing. We therefore build upon the coupling flow architecture originally proposed in Ref. 18 and refined in Refs. 19, 21 which allows to model ordered and disordered systems, such as solids and liquids, with the same architecture. Concretely, periodic boundary conditions can be respected utilizing circular splines [30], bijective transformations defined on a torus, whereas invariance under particle permutation can be achieved using (by construction) permutation-equivariant transformers [31] (see Sec. 4.3 for details).
Figure 2: Training results for 180 Lennard-Jones particles in FCC phase. a, KL divergence (Eq. (12)) between generated and target distributions at the evaluation state ${\bf c}_{\rm eval}$ obtained from a flow trained only on the evaluation state (orange) and from a flow trained in a conditional way (blue). Thermodynamic states ($T^{*}_{0},P^{*}_{0}$) and ($T^{*}_{\rm eval},P^{*}_{\rm eval}$) are marked as blue cross and blue star, respectively, in the middle panel. Learning rate reductions were applied for the conditional flow after 30 and 45 epochs. The orange dashed lines marks the training time after which the sampling efficiency of the single-point flow was evaluated. b and c, Sampling efficiency (Eq. (11)) for conditional and single-point flow, respectively, in % evaluated over the range of thermodynamic states used for training the conditional flow. d, Deviation in $\Delta G^{*}$ (relative to ($T^{*}_{0},P^{*}_{0}$)) of the conditional flow from MD+MBAR along the path defined by the black arrow in b. The dashed lines indicate one standard deviation evaluated over 10 flow runs, the shaded area correspond to the maximum standard deviation of 10 MD+MBAR runs. e and f, Radial distribution function of configurations from the prior, generated, and reweighted distributions of the conditional flow compared to MD at different thermodynamic states. g, 2D projections of 500 different configurations and mean box dimensions. The left-most panel contains samples from the prior distribution, the remaining panels show configurations generated for increasing temperature and decreasing pressure. For reference, prior configurations and mean box are superimposed to the generated configurations as orange shadows and dashed lines, respectively. In all plots, the fixed particle is located at (0,0).
To measure the performances of the two flow types over training time, we evaluate the KL divergence (Eq. (12), with $\Delta f$ from standard MD+MBAR simulations) between the generated distribution and the target distribution at the evaluation point (Fig. 2 a). The KL divergence of the single-point flow quickly converges within a few epochs, after which an increase in divergence indicates the onset of overfitting. The KL divergence of the conditional flow is much more noisy in the first 20 epochs of training, with a temporary increase before continuing to decrease. The convergence behavior can be explained by the fact that the conditional flow requires more training steps as it learns a whole family of mappings simultaneously. After applying a learning rate reduction, $D_{\text{KL}}$ becomes less noisy and converges to a similar value as the single-point flow. Even though training the conditional flow until convergence requires approximately 10 times more epochs than training the single-point flow, the additional training cost is amortized by the wide range of thermodynamic states covered by the conditional flow.
To assess the accuracy of the flow as free energy estimator, using 2 000 test samples, we compute the Gibbs free energy difference $\Delta G^{*}$ (Eq. (18)) with respect to the prior along the path $P^{*}\in[4,20]$ for a fixed $T^{*}=1.0$ and evaluate the error compared to results obtained with MBAR used across a set of MD simulations (Fig. 2 d). Overall, we observe that both methods are in excellent agreement over the entire range. The accuracy only decreases slightly with increasing distance from the prior and decreasing sampling efficiency. Nevertheless, even for the furthest point, $P^{*}=4$, the mean deviation per particle is only $5\cdot 10^{-3}$, while the absolute value in $\Delta G^{*}$ between $P^{*}=4$ and $P^{*}=20$ is around 14.5 per particle, that is the maximum relative error along the line is less than 0.05%. This demonstrates that the conditional flow can be used for the efficient evaluation of free energy differences also for states far away from the prior.
Furthermore, we illustrate the ability of the model to accurately capture pairwise correlations over varying thermodynamic conditions by computing the radial distribution function (RDF) $g(r)$ [5] for configurations at the start and end point of the path (Fig. 2 e and f) and compare to MD. For both $(T,P)$-conditions, the RDFs resulting from samples of the generated distributions are nearly indistinguishable from the MD results, even before reweighting, while samples from the prior distribution clearly yield different RDFs. Any remaining deviations between the mapped and MD results are resolved by reweighting the generated samples (Eq. (2)). The ability of the conditional flow model to generate samples with varying displacements and volumes is further illustrated in Fig. 2 g, where 2D-projections of 500 generated configurations with increasing temperature and decreasing pressure are shown.
2.3 Mapping out coexistence lines
Figure 3: Coexistence line a, Contourplot of $\Delta G^{*}=G^{*}_{\rm sol}-G^{*}_{\rm liq}$ per particle as obtained from the conditional flows. Liquid-solid coexistence corresponds to $\Delta G^{*}\!=\!0$. The dashed line denotes the coexistence line obtained from a high-accuracy MD+MBAR run. The locations of solid and liquid priors are marked as blue and red circles, respectively. b, Deviation in the melting temperature of the conditional flow from MD+MBAR. The uncertainties of the flow predictions were obtained as the standard deviation over 10 independent runs. The shaded area indicates the maximum fluctuations over 10 MD+MBAR runs.
Having demonstrated the efficiency of the conditional flow for accurately computing free energy differences over a wide range of thermodynamic states, we now apply this methodology to the exploration of the phase diagram of the LJ system.
Using the LJ system as an example, we demonstrate that our framework based on $(T,P)$-conditioned flows can alleviate these issues. Specifically, we compute the coexistence line between the solid FCC and liquid phase. The only user-defined input to our workflow consists in choosing appropriate locations of the prior states for the two phases of interest. If a point of coexistence is already known, this point is a good choice as the sampling efficiency is highest close to the prior location (see Fig. 2) which allows to compute the coexistence line with only a small number of samples. To emphasize that our approach works without explicit knowledge of the coexistence region, we define a broad range of thermodynamic states with $T^{*}\in[0.6,1.6]$ and $P^{*}\in[1,20]$ in which the coexistence line is expected [34]. Based on physical intuition, we place the prior for the solid at low temperature and high pressure $(T^{*}_{\rm sol}=1,P^{*}_{\rm sol}=14)$, and for the liquid at high temperature and low pressure $(T^{*}_{\rm liq}=1.4,P^{*}_{\rm liq}=6)$. Conditional flows for both phases are trained over the expected region of coexistence (details are given in the SI). After training, we define a grid of points in the $(T^{*},P^{*})$ diagram with a spacing of $(\Delta T^{*},\Delta P^{*})$=(0.05,1.4) and use the flows to evaluate the relative free energies to the respective prior based on 2000 samples.
In Fig. 3 a, the free energy difference between the solid and liquid phase around the coexistence line is shown as a function of temperature and pressure (see SI for free energy differences over the entire $(T,P)$-range considered during training). The coexistence line predicted by the flow is in excellent agreement with the reference values, obtained from high-accuracy MD+MBAR simulations (dashed line in Fig. 3 a), with only small deviations in the high-temperature region. In Fig. 3 b, the error in the melting temperatures $T^{*}_{\rm melt}$ as a function of pressure are shown as determined by the flow in comparison to MD+MBAR. For the low-pressure regions, the melting temperatures predicted by the flow are well within the error bars of the reference calculations. Only for higher pressures, deviations become slightly larger. Overall, the point of equal free energy, and therefore the melting temperatures, can be determined with a maximum (mean) deviation accuracy of less than $0.01\,T^{*}$.
2.4 Computational efficiency
[…]
3 Discussion
The introduced framework based on conditional normalizing flows provides a novel approach to efficiently calculate differences in free energies over a wide range of thermodynamic states. Contrary to traditional approaches, such as thermodynamic integration or expanded ensembles, it does not require an explicit discretization of the temperature and pressure space. For both solid and liquid phases, flows trained in a conditional fashion achieve large sampling efficiencies even far away from the prior states. As a consequence, only minimal prior knowledge about the system and the melting region is required to map out coexistence lines between two phases with high accuracy. Our workflow, therefore, constitutes a first step towards the efficient and simple calculation of phase diagrams using normalizing flows.
4 Methods
@@ 4.1 Traditional free energy estimators
We consider the isothermal-isobaric ($NPT$) ensembles, for which the reduced potential of a configuration ${\bf x}$ is defined as [4]
@@
From the partition function, the dimensionless configurational free energy $f=\beta G$ can be obtained as
@@
Since evaluating the integral over configuration space is unfeasible for most interaction potentials, evaluating the absolute free energy is usually impossible. A more tractable approach is to evaluate the difference in free energy between two states $A$ and $B$, $\Delta f_{AB}=f_{B}-f_{A}$, as
$\log\Delta f_{AB}=-\frac{Z_{B}}{Z_{A}}=\mathbb{E}_{({\bf x},V)\sim q_{A}}\biggl{[}\exp(-\Delta u_{AB}({\bf x},V))\biggr{]}$ (6)
with the difference in the reduced potential $\Delta u_{AB}({\bf x},V)=u_{B}({\bf x},V)-u_{A}({\bf x},V)$. Eq. (6) is known as Zwanzig identity [3], which is a central result of Free Energy Perturbation (FEP) and constitutes the foundation of many free energy estimators. While being formally exact, the convergence of FEP crucially depends on the overlap between the corresponding configurational distributions $q_{A}({\bf x},V)$ and $q_{B}({\bf x},V)$, making it a highly biased and noisy estimator in practical calculations, even for small molecules [39].
@@
Targeted free energy perturbation (TFEP) [6] tackles the problem of having intermediate states by defining an invertible map $\mathcal{M}$ transforming samples $({\bf x},V)$ from $q_{A}$ to samples $({\bf x}^{\prime},V^{\prime})=\mathcal{M}({\bf x},V)$ from a new target distribution $q^{\prime}_{A}$ which shares a larger overlap with $q_{B}$ than $q_{A}$. For a perfect map, $q^{\prime}_{A}=q_{B}$. Similarly, the inverse map $\mathcal{M}^{-1}$ transforms samples from $q_{B}$ to $q_{B}^{\prime}$ and $B^{\prime}=A$ for a perfect map. Using the Jacobian of the mapping, $J_{\mathcal{M}}({\bf x},V)$, the generated distribution can be computed analytically using the change of variable theorem
$q^{\prime}_{A}({\bf x}^{\prime},V^{\prime})=q_{A}({\bf x},V^{\prime})|\det J_{\mathcal{M}}({\bf x},V^{\prime})|^{-1}.$ (7)
@@
The key idea of learned free energy perturbation (LFEP) consists in parametrizing $\mathcal{M}$ as a normalizing flow which removes the need of crafting a suitable mapping based on physical intuition only [18, 41].
4.2 Conditioning normalizing flows on temperature and pressure
Normalizing flows are a class of deep generative models which are able to generate samples from a target distribution [14, 42]. Normalizing flows aim to learn the optimal set of parameters $\theta^{*}$ of an invertible function $f({\bf x})$ that maps samples ${\bf x}$ from a prior distribution $q_{A}({\bf x})$ into samples from a target distribution $q_{B}({\bf x})$. In practice, the learned set of parameters $\theta$ will be different from $\theta^{*}$ and the flow is trained to minimize the Kullback-Leibler (KL) divergence between the generated distribution $q_{A}^{\prime}({\bf x})$ and $q_{B}({\bf x})$. For physical systems with an equilbirium distribution defined in Eq. (4), the KL divergence can be expressed as [13, 18]:

### v2

@@ Abstract
The accurate prediction of phase diagrams is of central importance for both the fundamental understanding of materials as well as for technological applications in material sciences. However, the computational prediction of the relative stability between phases based on their free energy is a daunting task, as traditional free energy estimators require a large amount of simulation data to obtain uncorrelated equilibrium samples over a grid of thermodynamic states. In this work, we develop deep generative machine learning models based on the Boltzmann Generator approach for entire phase diagrams, employing normalizing flows conditioned on the thermodynamic states, e.g., temperature and pressure, that they map to. By training a single normalizing flow to transform the equilibrium distribution sampled at only one reference thermodynamic state to a wide range of target temperatures and pressures, we can efficiently generate equilibrium samples across the entire phase diagram. Using a permutation-equivariant architecture allows us, thereby, to treat solid and liquid phases on the same footing. We demonstrate our approach by predicting the solid-liquid coexistence line for a Lennard-Jones system in excellent agreement with state-of-the-art free energy methods while significantly reducing the number of energy evaluations needed.
  • August 16, 2024
@@ 1 Introduction
In recent years, machine learning (ML) models have found broad use in molecular simulations. ML interatomic potentials have been shown to strongly reduce the cost of ab initio MD simulations while retaining an accuracy comparable to density functional theory or other quantum chemistry methods [8, 9, 10]. Lately, these potentials have also been applied in the calculation of phase diagrams [11, 12] providing faster energy evaluations while the sampling challenge remains. Complementary, deep generative models are being developed to generate uncorrelated samples of molecular structures in one go, with the aim of overcoming the sampling problem typical of MD and MC. In particular, the Boltzmann Generator (BG) approach [13] leverages a combination of highly expressive invertible coordinate transformations, such as normalizing flows [14, 15], and statistical mechanics methods such as importance sampling or Metropolis-Hastings Monte Carlo, to generate independent identically distributed samples of the targeted equilibrium distribution. This approach has emerged both in the molecular sciences [13] as well as in lattice field theory [16, 17]. Within the TFEP-based learned free energy perturbation (LFEP), using flows enables to learn flexible mappings for systems where physical intuition is insufficient [18]. This approach has successfully been applied to sample small solvation systems [19], monoatomic solids [20, 21, 22], and small molecules with complex quantum-mechanical potentials based on cheaper potentials [23] or at different temperatures to overcome slow modes [24]. However, while LFEP offers an elegant way to learn a mapping between two specific thermodynamic states, it cannot be used efficiently for the calculation of phase diagrams since a new flow model needs to be trained for each thermodynamic state.
@@
In this work, we generalize the LFEP and BG approaches by employing normalizing flows conditioned on the target thermodynamic states. This conditional BG is capable of mapping MD samples from only one reference thermodynamic state across the entire $(T,P)$-range of the phase diagram with a single trained model. By training to match the equilibrium distributions at all thermodynamic states and appropriately reweighting the generated samples, free energy differences can be very efficiently computed for any temperature and pressure. In this way, the coexistence line between phases can be mapped out with only minimal prior knowledge of its location, while avoiding the need to conduct numerous MD simulations across a grid of thermodynamic states. Using a permutation-equivariant architecture [18], our approach further treats ordered and disordered phases on the same footing. We demonstrate our developed framework by determining the coexistence line between solid and liquid phases of a Lennard-Jones (LJ) system consisting of 180 interacting particles, achieving highly accurate free energy estimates for both phases over the entire range of temperatures and pressures.
2 Theoretical background
@@ 2.1 Traditional free energy estimators
We consider the isothermal-isobaric ($NPT$) ensemble, for which the reduced potential of a configuration ${\bf x}$ is defined as [4]
@@
From the partition function, the dimensionless configurational Gibbs free energy $f=\beta G$ can be obtained as
@@
Since evaluating the integral over configuration space is unfeasible for most interaction potentials, evaluating the absolute free energy is usually impossible. A more tractable approach consists in evaluating the difference in free energy between two states $A$ and $B$, $\Delta f_{AB}=f_{B}-f_{A}$. In particular, many of the most widely used free energy estimators are based on the Zwanzig identity [3], the central result of Free Energy Perturbation (FEP):
$\Delta f_{AB}=-\log\mathbb{E}_{({\bf x},V)\sim q_{A}}\biggl{[}\exp(-\Delta u_{AB}({\bf x},V))\biggr{]}$ (4)
Here, $\Delta u_{AB}({\bf x},V)=u_{B}({\bf x},V)-u_{A}({\bf x},V)$ is the difference in the reduced potential between the two states. While being formally exact, the convergence of FEP crucially depends on the overlap between the corresponding configurational distributions $q_{A}({\bf x},V)$ and $q_{B}({\bf x},V)$, making it a highly biased and noisy estimator in practical calculations, even for small molecules [25].
@@
Targeted free energy perturbation (TFEP) [6] tackles the problem of having intermediate states by defining an invertible map $\mathcal{M}$ transforming samples $({\bf x},V)$ from $q_{A}$ to samples $({\bf x}^{\prime},V^{\prime})=\mathcal{M}({\bf x},V)$ from a new target distribution $q^{\prime}_{A}$ which shares a larger overlap with $q_{B}$ than $q_{A}$. For a perfect map, $q^{\prime}_{A}=q_{B}$. Similarly, the inverse map $\mathcal{M}^{-1}$ transforms samples from $q_{B}$ to $q_{B}^{\prime}$ and $q^{\prime}_{B}=q_{A}$ for a perfect map. Using the Jacobian of the mapping, $J_{\mathcal{M}}({\bf x},V)$, the generated distribution can be computed analytically using the change of variable theorem
$q^{\prime}_{A}({\bf x}^{\prime},V^{\prime})=q_{A}({\bf x},V)|\det J_{\mathcal{M}}({\bf x},V)|^{-1}.$ (5)
@@
The key idea of learned free energy perturbation (LFEP) [18] consists in parametrizing $\mathcal{M}$ as a normalizing flow which removes the need of crafting a suitable mapping based on physical intuition only.
2.2 Normalizing flows
Normalizing flows are a class of invertible deep generative models which are able to generate samples from a target distribution [14, 28]. In contrast to other generative models, the use of invertible transformations allows to evaluate the generated distribution using the change-of-variables technique, making them an attractive sampling tool for the physical sciences. Normalizing flows aim to learn the optimal set of parameters $\theta^{*}$ of an invertible function $f({\bf x})$ that maps samples ${\bf x}$ from a prior distribution $q_{A}({\bf x})$ into samples from a target distribution $q_{B}({\bf x})$. In practice, the learned set of parameters $\theta$ will be different from $\theta^{*}$ and the flow is trained by minimizing the Kullback-Leibler (KL) divergence between the generated distribution $q_{A}^{\prime}({\bf x})$ and $q_{B}({\bf x})$.
BGs target equilbirium distributions similar to Eq. (2). In this case, the KL divergence can be expressed as [13, 18]: